Given that P = (-4, 5) and Q = (-6, 4), find the component form and magnitude of the vector PQ.



1
Expert's answer
2021-05-10T16:44:02-0400

Given that "P = (-4, 5)" and "Q = (-6, 4)," the component form of the vector "PQ" is "(-6-(-4), 4-5) = (-2, -1)" and magnitude is "\sqrt{(-2)^2+(-1)^2}=\sqrt{5}."